# The Strong Cosmic Censorship Conjecture

### v1

@@ 2.3 The Einstein equations of spacetime dynamics and their initial value problem
Returning to Laplacian determinism: inextendibility of the MGHD. The MGHD $(\mathcal{M},g)$ constructed in Theorem 2.4 consists of worldlines of inextendible timelike curves (i.e., observers). But what if the MGHD $(\mathcal{M},g)$ itself was extendible, i.e., there exists a Lorentzian manifold (¯$(\bar{\mathcal{M}},\bar{g)}$ solution of (2.4) and a proper embedding $J:\mathcal{M}\rightarrow\bar{\mathcal{M}}$? Due to the hyperbolic nature of (2.4), one cannot expect such an extension to be unique [26, 22], which already demonstrates a failure of global uniqueness for the initial value problem described above. This failure is also accompanied by a breakdown of Laplacian determinism, in the sense that the timelike curves which were inextendible in $(\mathcal{M},g)$ become extendible within (¯$(\bar{\mathcal{M}},\bar{g)}$: since the extension (¯$(\bar{\mathcal{M}},\bar{g)}$ is arbitrary, so is the fate of these observers, which is not uniquely determined by their initial conditions at the Cauchy surface $\Sigma$. The validity of Laplacian determinism thus mandates the inextendibility of the MGHD as a solution to the Einstein equations, which motivated our formulation of Strong Cosmic Censorship as Conjecture 1.1 in Section 1.4. To conclude this section, we will formulate three physically-relevant notions of MGHD inextendibility of decreasing stength:
@@ 3.2 The Schwarzschild black hole
The spacetime geometry of the Schwarzshild spacetime $(\mathbb{R}^{3+1},g_{S})$ is depicted in Figure 2, with the event [^9] The event horizon +=$\mathcal{H}^{+}=\{r=2M\}$ is a regular hypersurface for $g_{S}$ under the appropriate choice of coordinates [6, 31]. horizon +=$\mathcal{H}^{+}=\{r=2M\}$ and the spacelike singularity $\mathcal{S}=\{r=0\}$. The $\{r=0\}$ singularity is known to be terminal, in the sense that the Kretschmann scalar is infinite [46] (curvature blow-up): it is thus well-known that the Schwarzschild spacetime is $C^{2}$-inextendible [42] and, therefore, satisfies Strong Cosmic Censorship in the sense of Statement III in Section 2.3. Moreover, it is also classically known [46, 31] that in-falling observers reach $\mathcal{S}=\{r=0\}$ in finite proper time and experience infinite tidal deformations, which correspond, at least heuristically, to a metric blow-up on the $C^{0}$ level. These considerations beg the following question [42]: is the Schwarzschild spacetime inextendible in a stronger sense, e.g. as in Statement I of Section 2.3? The affirmative answer to this longstanding question was obtained by Sbierski [47] in the following theorem:
@@ 3.3 The Reissner–Nordström and Kerr black holes
Note that $r_{\pm}(M,e)$ correspond to the roots of $(1-\frac{2M}{r}+\frac{e^{2}}{r^{2}})$, with +=$\mathcal{H}^{+}=\{r=r_{+}(M,e)\}$ being the event horizon and $\mathcal{CH}^{+}=\{r=r_{-}(M,e)\}$ the Cauchy horizon: the MGHD of the Cauchy surface $\{t=0\}$ is $\{t\geq 0,\ r<r_{-}(M,e)\}$ and the Penrose diagram of the Reissner–Nordström spacetime for $0<|e|<M$ is the same as that of Figure 1. The extremal Reissner–Nordström spacetime corresponds to the case $|e|=M$. The rotating analogue of the Schwarzschild spacetime is the celebrated Kerr [48] black hole of angular momentum $0<|a|\leq M$ of the following form when $r\neq r_{\pm}(M,a)$:
@@
and is a solution of (2.4) in vacuum ($\mathbb{T}=0$). Analogously, +=$\mathcal{H}^{+}=\{r=r_{+}(M,a)\}$ is the event horizon and $\mathcal{CH}^{+}=\{r=r_{-}(M,a)\}$ the Cauchy horizon, where $r_{\pm}(M,a)=M\pm\sqrt{M^{2}-a^{2}}$; the MGHD is $\{t\geq 0,\ r<r_{-}(M,a)\}$ and the Penrose diagram of the Kerr spacetime for $0<|a|<M$ is given in Figure 1. Finally, the extremal Kerr spacetime corresponds to $|a|=M$.
@@ 4.3 Nonlinear dynamics and spherically-symmetric matter models
where $m^{2}>0$. The Einstein–Klein–Gordon equation, in contrast, is often viewed as a matter model in its own sake (as opposed to a toy model for vacuum dynamics) which features different phenomena than the Einstein vacuum equations [69, 70, 28, 71, 72].
@@ 4.4 Einstein dynamics: nonlinear stability of the Cauchy horizon
$[1+|v|^{s}]\phi_{|\mathcal{H}^{+}}(v)\text{ and its derivative are bounded in an unweighted-in-}v\text{ norm.}$. (4.6)
@@
[…]
@@ 4.7 Two-ended black holes with a spacelike singularity and a Cauchy horizon
Interior of charged hairy black holes. The author’s works [86] and [89] with Li considered the interior of black holes with a scalar hair– namely a non-trivial stationary scalar field on the event horizon. These spacetimes are spherically-symmetric solutions of the Einstein–Maxwell-scalar-field equations (4.2) in [86] and Einstein–Maxwell-charged-scalar-field (4.4) in [89] and their terminal boundary is entirely-spacelike, thus they do not possess any Cauchy horizon (compare with Theorem 4.4). They have a two-ended ×$\mathbb{R}\times\mathbb{S}^{2}$ spatial topology, but are not associated to an asymptotically flat [^15] However, Zheng recently [90] constructed a class of asymptotically AdS (the analogue of asymptotic flatness in the presence of a negative cosmological constant) black holes to which Theorem 4.11 applies, see the discussion in Section 7. black hole exterior, and thus do not play any direct role in Strong Cosmic Censorship as formulated in Conjecture 1.1. They will be used, however, in the asymptotically flat black hole constructions discussed in the paragraph below. For now, we state the results of [86, 89].
@@
There exists spatially-homogeneous solutions of the Einstein–Maxwell-(un)charged-scalar-field equations (4.2) or (4.4) with a constant scalar field data $\phi\equiv\epsilon\neq 0$ on a two-ended event horizon $\mathcal{H}_{L}\cup\mathcal{H}_{R}$, whose MGHD terminates at a spacelike boundary $\mathcal{S}=\{r=0\}$.
@@
    (violent collapse, [86]) For the Einstein–Maxwell-scalar-field equations (4.2), $\mathcal{S}=\{r=0\}$ is approximated by a Kasner metric of exponents $(1-2C\epsilon^{2},C\epsilon^{2},C\epsilon^{2})$ for a constant $C>0$.
@@
    (fluctuating collapse, [89]) For the Einstein–Maxwell-charged-scalar-field equations (4.4), $\mathcal{S}=\{r=0\}$ is approximated by a Kasner metric of positive exponents $(1-2p(\epsilon),p(\epsilon),p(\epsilon))$, where $p(\epsilon)\in(0,\frac{1}{2})$ is an algebraic function of a highly-oscillating quantity $\alpha(\epsilon)$ of the form
    $\alpha(\epsilon)\approx C\sin(\omega_{0}\epsilon^{-2}+O(\log(\epsilon^{-1})))\text{ as }\epsilon\rightarrow 0$
    for $C\neq 0$ and $\omega_{0}\in\mathbb{R}-\{0\}$. Moreover, for some values of $\epsilon$, a Kasner bounce takes place.
@@
If $(p_{1},p_{2},p_{3})=(1,0,0)$ (or perturbations), $g_{K}$ is flat (i.e., locally isometric to the Minkowski metric (3.1)) but otherwise, $\mathcal{S}=\{t=0\}$ is a spacelike singularity. In Theorem 4.11, the Kasner exponents $p_{i}$ are all positive (which is consistent with the known stability results in this case [92], see Section 7). Note, however, that in the uncharged-scalar-field case (4.2), the Kasner exponents degenerate to the Minkowskian endpoint $(p_{1},p_{2},p_{3})=(1,0,0)$ when $\epsilon$ becomes arbitrarily small.
@@
The terminology “Kasner bounce” originates from the foundational works of BKL [93, 94] on Kasner dynamics for the Einstein equations. These celebrated heuristics proposed an instability mechanism for Kasner metrics (4.10) with one negative $p_{i}$ leading to the transition (“Kasner epoch”) to a stable Kasner metric with all positive $p_{i}$’s. Statement 2 of Theorem 4.11 provides the first rigorous quantitative study of the long-hypothesized Kasner bounce phenomenon.
@@
Construction of asymptotically flat two-ended black holes. A gluing argument [^16] While the gluing argument of [89] only provides two-ended asymptotically-flat black holes featuring a piece of hairy black hole singularity $\mathcal{S}$, we conjecture that one-ended versions of such construction are also possible. Unlike their two-ended counterparts, one-ended black holes play a major role in gravitational collapse, see Section 5. allows to construct two-ended asymptotically flat black holes satisfying the assumptions of Theorem 4.4, but whose terminal boundary features a spacelike component $\mathcal{S}^{\prime}$ isometric to a subset of $\mathcal{S}=\{r=0\}$ in Theorem 4.11 (see [89][Section 1.2] for details). In particular, one can in principle construct spherically-symmetric asymptotically flat solutions of Einstein–Maxwell-charged-scalar-field equations (4.4) exhibiting the Kasner bounce described in the previous paragraph.
In the case of the Einstein–Maxwell-charged-scalar-field equations (4.4), however, one would additionally require decay estimates in the black hole exterior to finalize the construction (see the discussion in Section 6). For the Einstein–Maxwell-scalar-field equations (4.2) case, these estimates have been obtained in [74, 84] and thus the gluing argument provides a large class of two-ended asymptotically flat black holes with the rightmost Penrose diagram of Figure 7, where part of the singularity $\mathcal{S}=\{r=0\}$ is spacelike and described by the Kasner-like asymptotics of Theorem 4.11.
We note that for such constructions, Theorem 4.11 only provides estimates near $\mathcal{S}$ away from the intersection $\mathcal{CH}^{+}\cap\mathcal{S}$. However, the results of [95] discussed in the next section apply to this situation and provide a quantitative description of the null-spacelike transition region near $\mathcal{CH}^{+}\cap\mathcal{S}$, see in particular Section 5.4.
@@ 5.1 Mathematical setting of gravitational collapse
Gravitational collapse is the widely-accepted theory accounting for the existence of astrophysical black holes, as pioneered in the celebrated work of Oppenheimer–Snyder [49] (recall the discussion in Section 3.4). In this astrophysical process, a star whose initial state is free of any trapped surface develops a trapped surface at a later time. The mathematical setting for gravitational collapse imposes [^17] Topological censorship considerations regarding an initial hypersurface $\Sigma$ indeed indicate that $\Sigma$ must have a trivial topology (like $\mathbb{R}^{3}$ in Definition 5.1) if it is free of any trapped surface, see [96] for interesting results in this direction. an initial surface $\Sigma$ that is diffeomorphic to $\mathbb{R}^{3}$ and free of any trapped surface, with the origin of $\mathbb{R}^{3}$ corresponding to the center of the star $\Gamma$. Of course, not all such spacetimes undergo a gravitational collapse process (e.g. recall from Section 3.1 that perturbations of the Minkowski spacetime do not admit any trapped surface). Nonetheless, in Definition 5.1 below, we introduce spacetimes that are initially free of any trapped surface: thus gravitational collapse may take place if the conditions leading to the formation of a trapped surface [^18] There is a rich mathematical theory for the formation of trapped surface, see e.g. [24, 66, 97, 98, 99, 100, 101]. are satisfied at a later time. For simplicity, we restrict ourselves to the vacuum case ((2.4) with $\mathbb{T}=0$), keeping in mind that Definition 5.1 generalizes in the presence of a suitable matter model (see e.g. [13]).
@@
Note that a spacetime suitable for gravitational collapse possesses one asymptotically flat end (“one-ended spacetime”), as opposed to the two-ended spacetimes (including the Reissner–Nordström (3.3) and Kerr (3.5) MGHDs) discussed in Section 4.6 whose topology is that of ×$\mathbb{R}\times\mathbb{S}^{2}$.

### v2

@@ 2.3 The Einstein equations of spacetime dynamics and their initial value problem
Returning to Laplacian determinism: inextendibility of the MGHD. The MGHD $(\mathcal{M},g)$ constructed in Theorem 2.4 consists of worldlines of inextendible timelike curves (i.e., observers). But what if the MGHD $(\mathcal{M},g)$ itself was extendible, i.e., there exists a Lorentzian manifold $(\bar{\mathcal{M}},\bar{g)}$ solution of (2.4) and a proper embedding $J:\mathcal{M}\rightarrow\bar{\mathcal{M}}$? Due to the hyperbolic nature of (2.4), one cannot expect such an extension to be unique [26, 22], which already demonstrates a failure of global uniqueness for the initial value problem described above. This failure is also accompanied by a breakdown of Laplacian determinism, in the sense that the timelike curves which were inextendible in $(\mathcal{M},g)$ become extendible within $(\bar{\mathcal{M}},\bar{g)}$: since the extension $(\bar{\mathcal{M}},\bar{g)}$ is arbitrary, so is the fate of these observers, which is not uniquely determined by their initial conditions at the Cauchy surface $\Sigma$. The validity of Laplacian determinism thus mandates the inextendibility of the MGHD as a solution to the Einstein equations, which motivated our formulation of Strong Cosmic Censorship as Conjecture 1.1 in Section 1.4. To conclude this section, we will formulate three physically-relevant notions of MGHD inextendibility of decreasing stength:
@@ 3.2 The Schwarzschild black hole
The spacetime geometry of the Schwarzshild spacetime $(\mathbb{R}^{3+1},g_{S})$ is depicted in Figure 2, with the event [^9] The event horizon $\mathcal{H}^{+}=\{r=2M\}$ is a regular hypersurface for $g_{S}$ under the appropriate choice of coordinates [6, 31]. horizon $\mathcal{H}^{+}=\{r=2M\}$ and the spacelike singularity $\mathcal{S}=\{r=0\}$. The $\{r=0\}$ singularity is known to be terminal, in the sense that the Kretschmann scalar is infinite [46] (curvature blow-up): it is thus well-known that the Schwarzschild spacetime is $C^{2}$-inextendible [42] and, therefore, satisfies Strong Cosmic Censorship in the sense of Statement III in Section 2.3. Moreover, it is also classically known [46, 31] that in-falling observers reach $\mathcal{S}=\{r=0\}$ in finite proper time and experience infinite tidal deformations, which correspond, at least heuristically, to a metric blow-up on the $C^{0}$ level. These considerations beg the following question [42]: is the Schwarzschild spacetime inextendible in a stronger sense, e.g. as in Statement I of Section 2.3? The affirmative answer to this longstanding question was obtained by Sbierski [47] in the following theorem:
@@ 3.3 The Reissner–Nordström and Kerr black holes
Note that $r_{\pm}(M,e)$ correspond to the roots of $(1-\frac{2M}{r}+\frac{e^{2}}{r^{2}})$, with $\mathcal{H}^{+}=\{r=r_{+}(M,e)\}$ being the event horizon and $\mathcal{CH}^{+}=\{r=r_{-}(M,e)\}$ the Cauchy horizon: the MGHD of the Cauchy surface $\{t=0\}$ is $\{t\geq 0,\ r<r_{-}(M,e)\}$ and the Penrose diagram of the Reissner–Nordström spacetime for $0<|e|<M$ is the same as that of Figure 1. The extremal Reissner–Nordström spacetime corresponds to the case $|e|=M$. The rotating analogue of the Schwarzschild spacetime is the celebrated Kerr [48] black hole of angular momentum $0<|a|\leq M$ of the following form when $r\neq r_{\pm}(M,a)$:
@@
and is a solution of (2.4) in vacuum ($\mathbb{T}=0$). Analogously, $\mathcal{H}^{+}=\{r=r_{+}(M,a)\}$ is the event horizon and $\mathcal{CH}^{+}=\{r=r_{-}(M,a)\}$ the Cauchy horizon, where $r_{\pm}(M,a)=M\pm\sqrt{M^{2}-a^{2}}$; the MGHD is $\{t\geq 0,\ r<r_{-}(M,a)\}$ and the Penrose diagram of the Kerr spacetime for $0<|a|<M$ is given in Figure 1. Finally, the extremal Kerr spacetime corresponds to $|a|=M$.
@@ 4.3 Nonlinear dynamics and spherically-symmetric matter models
and $m^{2}>0$. For instance, for an uncharged scalar field, we have the Einstein–Maxwell–Klein–Gordon equations:
$\mathbb{T}_{\mu\nu}=\mathbb{T}_{\mu\nu}^{EM}+\mathbb{T}_{\mu\nu}^{KG},\ \Box_{g}\phi=m^{2}\phi,\ \nabla^{\mu}F_{\mu\nu}=0,\ \mathbb{T}_{\mu\nu}^{KG}=\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}[g^{\alpha\beta}\partial_{\alpha}\phi\partial_{\beta}\phi+m^{2}|\phi|^{2}]g_{\mu\nu}.$ (4.6)
The Einstein–(Maxwell)–Klein–Gordon equation, in contrast to the previously-mentioned models, is often viewed as a matter model in its own sake (as opposed to a toy model for vacuum dynamics) which features different phenomena than the Einstein vacuum equations [69, 70, 28, 71, 72].
@@ 4.4 Einstein dynamics: nonlinear stability of the Cauchy horizon
[…]
@@
$\bigl|\int_{1}^{\infty}\phi_{|\mathcal{H}^{+}}(v)e^{-i[\omega_{res}+o(1)]v}dv\bigr|<\infty,$ (4.9)
@@ 4.7 Two-ended black holes with a spacelike singularity and a Cauchy horizon
Interior of charged hairy black holes. The author’s works [86] and [89] with Li considered the interior of black holes with a scalar hair– namely a non-trivial stationary scalar field on the event horizon. These spacetimes are spherically-symmetric solutions of the Einstein–Maxwell–Klein–Gordon equations (4.6) in [86] and Einstein–Maxwell-charged-scalar-field (4.4) in [89] and their terminal boundary is entirely-spacelike, thus they do not possess any Cauchy horizon (compare with Theorem 4.4). They have a two-ended $\mathbb{R}\times\mathbb{S}^{2}$ spatial topology, but are not associated to an asymptotically flat [^15] However, Zheng recently [90] constructed a class of asymptotically AdS (the analogue of asymptotic flatness in the presence of a negative cosmological constant) black holes to which Theorem 4.11 applies, see the discussion in Section 7. black hole exterior, and thus do not play any direct role in Strong Cosmic Censorship as formulated in Conjecture 1.1. They will be used, however, in the asymptotically flat black hole constructions discussed below. For now, we state the results of [86, 89].
@@
There exists spatially-homogeneous solutions of the Einstein–Maxwell equations coupled with a charged and/or massive scalar field, which is initially constant $\phi\equiv\epsilon\neq 0$ on a two-ended event horizon $\mathcal{H}_{L}\cup\mathcal{H}_{R}$ and obeying the following properties.
@@
    (violent collapse, [86]). For the Einstein–Maxwell-Klein–Gordon (4.6) equations, there exists a discrete set $\mathcal{C}$ with $0\in\mathcal{C}$, such that, if $m^{2}\in\mathbb{R}-\mathcal{C}$, the MGHD terminates at a spacelike boundary $\mathcal{S}=\{r=0\}$ and $\mathcal{S}=\{r=0\}$ is approximated by a Kasner metric of exponents $(1-2C\epsilon^{2},C\epsilon^{2},C\epsilon^{2})$ for some $C>0$.
@@
    (fluctuating collapse, [89]). For the Einstein–Maxwell-charged-scalar-field equations (4.4), if $q_{0}\neq 0$, then the MGHD terminates at a spacelike boundary $\mathcal{S}=\{r=0\}$ and $\mathcal{S}=\{r=0\}$ is approximated by a Kasner metric of positive exponents $(1-2p(\epsilon),p(\epsilon),p(\epsilon))$, where $p(\epsilon)\in(0,\frac{1}{2})$ is an algebraic function of a highly-oscillating quantity $\alpha(\epsilon)$ of the form α(ϵ) C sin(ω_0 ϵ^-2 +O(log(ϵ^-1))) as ϵ0 for $C\neq 0$ and $\omega_{0}\in\mathbb{R}-\{0\}$. Moreover, for some values of $\epsilon$, a Kasner bounce takes place.
@@
If $(p_{1},p_{2},p_{3})=(1,0,0)$ (or perturbations), $g_{K}$ is flat (i.e., locally isometric to the Minkowski metric (3.1)) but otherwise, $\mathcal{S}=\{t=0\}$ is a spacelike singularity. In Theorem 4.11, the Kasner exponents $p_{i}$ are all positive (which is consistent with the known stability results in this case [92], see Section 7). Note, however, that in the uncharged-scalar-field case of [86], the Kasner exponents degenerate to the Minkowskian endpoint $(p_{1},p_{2},p_{3})=(1,0,0)$ when $\epsilon$ becomes arbitrarily small.
@@
The terminology “Kasner bounce” originates from the foundational works of BKL [93, 94] on Kasner dynamics for the Einstein equations. These celebrated heuristics proposed an instability mechanism for Kasner metrics (4.11) with one negative $p_{i}$ in three space-dimensions, leading to the transition (“Kasner epoch”) to a stable Kasner metric with all positive $p_{i}$’s. Statement 2 of Theorem 4.11 provides the first rigorous quantitative study of the long-hypothesized Kasner bounce phenomenon; see also the recent works [95, 96] for the first examples of such Kasner bounces without assuming spatial homogeneity.
@@
Construction of asymptotically flat two-ended black holes. A gluing argument allows to construct two-ended asymptotically flat black holes satisfying the assumptions of Theorem 4.4 and thus with a non-empty Cauchy horizon $\mathcal{CH}^{+}$, but whose terminal boundary also features a spacelike component $\mathcal{S}^{\prime}$ isometric to a subset of $\mathcal{S}=\{r=0\}$ in Theorem 4.11 (see [89][Section 1.2] for details). In particular, one can in principle construct spherically-symmetric asymptotically flat solutions of Einstein–Maxwell-charged-scalar-field equations (4.4) exhibiting the Kasner bounce described in the previous paragraph. However, one would additionally require decay estimates in the black hole exterior–which are open at present–to finalize the construction of two-ended asymptotically flat black holes featuring a null Cauchy horizon and a spacelike singularity; this is true both in the massive case of [86] ((4.6)) and the charged case of [89] ((4.4)), see the related discussions in Section 6.
Independently of the constructions of Theorem 4.11, however, the author proved [97] that there exist two-ended asymptotically flat black holes solutions of Einstein–Maxwell-(uncharged)-scalar-field equations (4.2) as in the right panel of Figure 7, complementing Theorem 4.7 in showing both possibilities of Figure 7 are possible.
Theorem 4.12 ([97]).
There exists a large class of spherically symmetric two-ended asymptotically flat black hole solution of (4.2) such that the terminal boundary has two non-empty components: a weakly singular null Cauchy horizon $\mathcal{CH}^{+}\neq$$\emptyset$, and a crushing singularity $\mathcal{S}=\{r=0\}$. Moreover, $\mathcal{S}$ is spacelike near $\mathcal{CH}^{+}\cap\mathcal{S}$.
We have not yet addressed the spacetime behavior near the null-to-spacelike transition at $\mathcal{CH}^{+}\cap\mathcal{S}$ in Theorem 4.12. However, this transition is precisely characterized by the results of [98], which are presented in the following Section 5 in the context of gravitational collapse, see Theorem 5.4.
@@ 5.1 Mathematical setting of gravitational collapse
Gravitational collapse is the widely-accepted theory accounting for the existence of astrophysical black holes, as pioneered in the celebrated work of Oppenheimer–Snyder [49] (recall the discussion in Section 3.4). In this astrophysical process, a star whose initial state is free of any trapped surface develops a trapped surface at a later time. The mathematical setting for gravitational collapse imposes [^16] Topological censorship considerations regarding an initial hypersurface $\Sigma$ indeed indicate that $\Sigma$ must have a trivial topology (like $\mathbb{R}^{3}$ in Definition 5.1) if it is free of any trapped surface, see [99] for interesting results in this direction. an initial surface $\Sigma$ that is diffeomorphic to $\mathbb{R}^{3}$ and free of any trapped surface, with the origin of $\mathbb{R}^{3}$ corresponding to the center of the star $\Gamma$. Of course, not all such spacetimes undergo a gravitational collapse process (e.g. recall from Section 3.1 that perturbations of the Minkowski spacetime do not admit any trapped surface). Nonetheless, in Definition 5.1 below, we introduce spacetimes that are initially free of any trapped surface: thus gravitational collapse may take place if the conditions leading to the formation of a trapped surface [^17] There is a rich mathematical theory for the formation of trapped surface, see e.g. [24, 66, 100, 101, 102, 103]. are satisfied at a later time. For simplicity, we restrict ourselves to the vacuum case ((2.4) with $\mathbb{T}=0$), keeping in mind that Definition 5.1 generalizes in the presence of a suitable matter model (see e.g. [13]).
@@
Note that a spacetime suitable for gravitational collapse possesses one asymptotically flat end (“one-ended spacetime”), as opposed to the two-ended spacetimes (including the Reissner–Nordström (3.3) and Kerr (3.5) MGHDs) discussed in Section 4.6 whose topology is that of $\mathbb{R}\times\mathbb{S}^{2}$.